What is the slope of a line?
Given two points on a line, how is the slope calculated?
Given the equation of a line, how is the slope calculated?
Demonstrate by determining the slope of "3x + 5y = 19."

Answer by checkley77(12569)

Slope=(Y2-Y1)/(X2-X1) is the slope formula given 2 points.
Y=mX+b
m=slope.
Y=4x-5 4 is this lines slope.
3x+5y=19
5y=-3x+19
y=-3x/5+19/5
Slope=-3/5.